A population of birds contains 23 animals with red tail feathers and 66 animals with blue tail feathers. Blue tail feathers are the dominant trait. (Answer in percent or decimal)


A. What is the frequency of the red allele?


B. What is the frequency of the blue allele?


C. What is the frequency of heterozygotes?


D. What is the frequency of birds homozygous for the blue allele?



1
Expert's answer
2021-02-08T10:21:27-0500

The genotypes are:

AA (blue), Aa (blue), aa (red).

The frequency of red genotype is:

f(aa) = 23 / (23 + 66) = 0.26

A. The frequency of red allele is:

f(a) = a1/2 = 0.51

B. The frequency of blue allele is:

f (A) = 1 - f(a) = 1 - 0.51 = 0.49

C. The frequency of heterozygotes is:

f (Aa) = 2 × f (a) × f (A) = 2 × 0.51 × 0.49 = 0.5

D. The frequency of birds homozygous for the blue allele:

f (AA) = f (A)2 = 0.492 = 0.24